# Simplifying generic smooth maps to the $2$–sphere and to the plane

### v1

@@ 5. Maps into $\mathbb{R}^{2}$
In this section, we show the following.
@@
The above theorem holds also for $n=2$; in fact, for each closed connected surface $M$, we can explicitly construct a generic map as described in the theorem. When $M$ is orientable, this is straightforward. For the non-orientable case, we can first construct an image simple generic map into $S^{2}$ by using [27, Theorem 1.4 (2)], for example, and then compose it with an appropriate generic map $S^{2}\to\mathbb{R}^{2}$ to get a desired generic map.
@@
There exists a generic map $f:M\to S^{2}$ which is image simple by Theorem 4.3. Let $D_{S}\cup A\cup D_{N}=S^{2}$ be a decomposition of $S^{2}$ such that $D_{S}$ and $D_{N}$ are disjoint $2$–disks, that $A$ is an annulus, that they are attached along their boundaries, and that $\mathop{\mathrm{Int}}\nolimits{D_{N}}\supset f(S(f))$. For a regular value $q\in D_{S}$, set $F=f^{-1}(q)$, which is an $(n-2)$–dimensional closed manifold. Note that $f|_{f^{-1}(D_{S}\cup A)}$ is a trivial smooth $F$–bundle over $D_{S}\cup A$ by Ehresmann’s fibration theorem. In particular, we have $f^{-1}(A)\cong F\times[-1,1]\times S^{1}$.
Let $h:F\to[1,2]$ be a Morse function such that the critical values are all distinct. Define the Morse function $\tilde{h}:F\times[-1,1]\to[1,3]$ by $\tilde{h}(x,t)=h(x)\cos(\pi t/2)+1$, $(x,t)\in F\times[-1,1]$ (refer to [2, Fig. 40]). Then, we can construct a stable map $g:M\to\mathbb{R}^{2}$ in such a way that $g|_{f^{-1}(D_{N})}=i_{N}\circ f|_{f^{-1}(D_{N})}$, $g|_{f^{-1}(D_{S})}=i_{S}\circ f|_{f^{-1}(D_{S})}$, and $g|_{f^{-1}}(A)$ is the composition of a trivial $S^{1}$–family of Morse functions $\tilde{h}\times\mathrm{id}_{S^{1}}:(F\times[-1,1])\times S^{1}\to[1,3]\times S^{1}$ and an appropriate embedding $[1,3]\times S^{1}\to\mathbb{R}^{2}$, where $i_{N}:D_{N}\to S^{2}$ and $i_{S}:D_{S}\to S^{2}$ are appropriate embeddings (see Fig. 13). For details, refer to [2, Proof of Theorem 7.1].
@@ 6. Relation to open book structures
In this section, we given an application of the above results in the odd dimensional case to open book structures.
@@
For some historical account of open book structures, the author is referred to Quinn [14, §2].
@@
Let $M$ be a compact $n$–dimensional manifold with boundary. Suppose that $B$ is a proper $(n-2)$–dimensional compact submanifold of $M$ with trivial normal bundle. Here a submanifold $B$ is proper if $B\cap\partial M=\partial B$ and $B$ intersects $\partial M$ transversely. Let $\tau:N(B)\to B\times D^{2}$ be a trivialization of the normal disk bundle $N(B)$ of $B$ in $M$, where $N(B)$ is identified with a closed tubular neighborhood of $B$ in $M$ such that $N(B)\cap\partial M$ can be identified with the tubular neighborhood $N(\partial B)$ in $\partial M$. We assume that there exists a smooth $S^{1}$–bundle $\varphi:M\setminus B\to S^{1}$ such that $\varphi|_{N(B)\setminus B}$ coincides with the composition (6.1) as above and that $\varphi|_{\partial M\setminus\partial B}$ induces an open book structure with binding $\partial B$ on $\partial M$. Such a structure of $M$ is also called an open book structure; $B$ is called the binding and the closure of each fiber of $\varphi$ is called a page. Note that the restriction of the structure to $\partial M$ gives rise to an open book structure of $\partial M$. In this case, we say that the open book structure on the boundary extends through $M$.
As an application of our results, we give a new proof to the following result originally due to Quinn [14] (see also [10, 22, 25]).
@@
Let $bB\subset\partial M$ be the binding, $b\tau:N(bB)\to bB\times D^{2}$ the trivialization, and $b\varphi:\partial M\setminus bB\to S^{1}$ the fibration associated with the open book structure on $\partial M$. Let $f|_{N(bB)}:N(bB)\to\mathbb{R}^{2}$ be defined by the composition
@@
Note that the number of cusps of $f$ is even, since the Euler characteristic of the pair $(M,F)$ vanishes, where $F$ is a fiber of $f$ over a point in $\partial\Delta$. Then, by the same procedure as in Levine’s cusp elimination theorem [11], we can eliminate the cusps of $f$ by pairs, since $M$ is connected. So, we may assume that $f$ has no cusps.
Now, we apply Lemma 4.8, taking into account the comments in Remark 4.9 (1) and (2), first to the definite fold image. As a result, we get a generic map $M\to\mathbb{R}^{2}$, denoted by $f$ again, such that $f(M)$ coincides with $\Delta$, the definite fold image is outward directed, and that the other fold images lie inside the interior of a $2$–disk, say $\Delta_{1}$, contained in the innermost region. Then, we apply Lemma 4.8 to $f$ restricted to $f^{-1}(\Delta_{1})$ to get a new generic map $f$ such that the fold image of absolute index $1$ is contained in $\mathop{\mathrm{Int}}\nolimits\Delta_{1}$, that it is outward directed, and that the other fold images lie inside the interior of a $2$–disk, say $\Delta_{2}$, contained in the innermost region. Repeating this procedure, we get a generic map $f:M\to\mathbb{R}^{2}$ such that $f(M)=\Delta$, that the fold images of absolute indices $i$ with $0\leq i<\lfloor(n-1)/2\rfloor$ are all directed outward, that their positions are consistent with the order of absolute indices (smaller indices in the outer position), and that the fold image of absolute index $i_{0}=\lfloor(n-1)/2\rfloor$ lies in the interior of a $2$–disk $\Delta_{i_{0}}$ contained in the innermost region. We may continue to assume that $f$ satisfies the conditions (1) and (2) of Definition 2.3.
@@
Now, we apply the always-realizable moves as depicted in Figs. 1417. We may assume that the disks $\Delta_{0},\Delta_{1},\ldots,\Delta_{i_{0}}$ are concentric, i.e., their centers all coincide with the origin $\mathbf{0}\in\mathbb{R}^{2}$. We denote by $\pi:\Delta\setminus\{\mathbf{0}\}\to S^{1}$ the radial projection. Let $A$ be the annulus $\Delta_{i_{0}}\setminus\mathop{\mathrm{Int}}\nolimits{\Delta_{i_{0}+1}}$, where $\Delta_{i_{0}+1}$ is a concentric disk in $\mathop{\mathrm{Int}}\nolimits{\Delta_{i_{0}}}$. We may assume that the fold image of absolute index $i_{0}$ lies in $\mathop{\mathrm{Int}}\nolimits{A}$ and that $\pi\circ f|_{S_{i_{0}}}:S_{i_{0}}\to S^{1}$ is an $S^{1}$–valued Morse function with distinct critical values, where $S_{i_{0}}$ is the set of folds of absolute index $i_{0}$ of $f$. Then, for each of the critical points, we apply the type II moves as depicted in Fig. 14 so that the images of the critical points by $f$ lie inside $\mathop{\mathrm{Int}}\nolimits{\Delta_{i_{0}+1}}$. (Before doing so, we arrange $f(S_{i_{0}})$ so that all the crossings are pushed into the regions in $A$ outside of those regions in which $f(S_{i_{0}})$ is depicted in the figure.) Note that in our case, only the top type II move in Fig. 5 is allowed and that the bottom type II move is not. Note also that as $n$ is odd, even if we reverse the normal orientation for a fold image of absolute index $i_{0}=(n-1)/2$, the absolute index does not change.
@@
We orient $S^{1}=\partial\Delta$ counter clockwise, for the moment, then with respect to that orientation, the numbers of local minima and maxima of the Morse function $\pi\circ f|_{S_{i_{0}}}$ are the same. Then, by further applying the type II moves, we may assume that the images of the critical points by $f$ lie in pairs, “face-to-face” as in the first figures of Figs. 15 and 16. Here, we can show that the fibers of $f$ over the points in $\Delta_{i_{0}+1}$ are connected by the same argument as in the proof of Lemma 4.10. Then, we apply always-realizable moves as depicted in Fig. 15 for each pair of the images of a local minimum and a local maximum. More precisely, we first apply type I moves to create 4 cusps. Then, we apply the cusp merge moves to the 2 pairs of cusps, which are allowed, since the fibers are connected. If we reverse the normal orientation of the newly born index–$(i_{0}+1)$ circle, then it turns into a fold image circle of absolute index $i_{0}-1$. Then, by applying the moves as depicted in Fig. 10 with $i_{0}$ replaced by $i_{0}-1$, we can turn this circle into a circle of absolute index $i_{0}$. Then, by applying the type II move once, we reach the final configuration in Fig. 15.
@@
Figure 15. A series of moves for a pair of critical point images
@@
Let $\Delta_{i_{0}+2}$ be a small closed $2$–disk centered at $\mathbf{0}$ disjoint from $f(S(f))$. Then, as $f$ restricted to $f^{-1}(\Delta_{i_{0}+2})$ and to $f^{-1}(\Delta_{i_{0}+2})\cap\partial M$ are submersions, we see that $f|_{f^{-1}(\Delta_{i_{0}+2})}:f^{-1}(\Delta_{i_{0}+2})\to\Delta_{i_{0}+2}$ is a trivial bundle. Furthermore, setting $B=f^{-1}(\mathbf{0})$, we see that $\partial B=bB$ and that the composition
@@
is a smooth fiber bundle, since $\pi\circ f$ restricted to $S(f)$ is a submersion. Hence, this gives an open book structure for $M$. Making $N(bB)$ smaller, we see that an open book structure naturally isomorphic to the given one extends through $M$. This completes the proof. ∎
@@
Note that Theorem 6.3 also covers the case with $\partial M=\emptyset$: just ignore the boundary open book structures in the proof.
@@ 7. Elimination of indefinite folds of certain indices
For the extended open book structure on $M$ constructed in the above proof of Theorem 6.3, the binding $B$ is connected, and its fundamental group satisfies $\pi_{1}(M)\cong\pi_{1}(B)$ if $n\geq 7$.
@@
Let us decompose $\Delta$, the unit disk in $\mathbb{R}^{2}$, into two half disks as $\Delta=\Delta_{+}\cup\Delta_{-}$, where $\Delta_{+}=\{(x,y)\in\Delta\,|\,y\geq 0\}$ and $\Delta_{-}=\{(x,y)\in\Delta\,|\,y\leq 0\}$. Then, we see that the inclusions $F\to f^{-1}(\Delta_{\pm})$ are homotopy equivalences. Now, $X=f^{-1}(\Delta_{+})\cap f^{-1}(\Delta_{-})$ is diffeomorphic to the union of $F$ and its copy attached along $B$ and hence the inclusion $B\to X$ induces an isomorphism between the fundamental groups by the van Kampen theorem. Finally, by the van Kampen theorem applied to $M=f^{-1}(\Delta_{+})\cup_{X}f^{-1}(\Delta_{-})$, we see that the inclusion $B\to M$ induces an isomorphism between the fundamental groups. This completes the proof. ∎
@@
Let $M$ be a closed simply connected $n$–dimensional manifold with $n\geq 7$ odd. Then, there exists a stable map $f:M\to\mathbb{R}^{2}$ without cusp points such that the absolute indices $i$ of the folds are either equal to 00 or satisfy $22\leq i\leq(n-1)/$ 2.
@@
As in the proof of Lemma 7.1, we consider the Morse function $g=f|_{F}:F\to[0,1]$, where $F$ is a compact $(n-1)$–dimensional manifold with boundary $\partial F=B$. By the construction of $f$, we see that the indices $i$ of the critical points satisfy $(n-1)/2\leq i\leq n-1$. On the other hand, by our assumption, $B$ and $F$ are simply connected.
If $g$ has two or more critical points of index $n-1$, i.e. local maxima, then there exists a critical point of index $n-2$ in such a way that its dual handle connects two dual 00–handles, since $F$ is connected. Therefore, either of the two critical points of index $n-1$ and the index $n-2$ critical point make a canceling pair. (Note that by reversing the orientation of $J=[0,1]$, they make a canceling 001111 pair of critical points.) Then by [8, Part I, Chapter V, Proposition 1.1], we can cancel them leaving a pair of cusps of absolute index 00. By repeating this procedure, we may assume that $g$ has only one maximal point (refer to Fig. 18).
@@
Let $U$ be a small open regular neighborhood of $J$ in $\mathbb{R}^{2}$. Note that $\Delta\setminus U$ is naturally identified with $[0,1]\times[-1,1]$, where the first factor corresponds to a radial outward arc and the second factor corresponds to the direction of $\partial\Delta=S^{1}$. Then, $f$ restricted to $f^{-1}(\Delta\setminus U)$ is considered to be a smooth $1$–parameter family of functions to $[0,1]$ parametrized by $[-1,1]$. Now, by applying an argument similar to that in the proof of [8, Part I, Chapter V, Proposition 3.5], we can homotopically modify $f$ so that it has only one fold circle of absolute index 00 and that it is embedded into $\mathbb{R}^{2}$ in the outermost place with the expense of creating some fold circles of absolute index $2$.
Since $F$ is simply connected of dimension greater than or equal to $6$, by [13, Theorem 8.1], we can cancel each critical point of index $1$ with the expense of creating critical points of index $3$. Then, again by the same argument as in the proof of [8, Part I, Chapter V, Proposition 3.5], we can homotopically modify $f$ so that it has no fold circle of absolute index $1$ and that it has only one fold circle of absolute index 00 embedded into $\mathbb{R}^{2}$ in the outermost place. This completes the proof. ∎
@@
We may assume that $f$ is a stable map. Then, as $f$ does not have folds of absolute index $1$, we can show that $W_{f}$ has the structure of a compact connected smooth $2$–dimensional manifold (for example, see [26, Theorem 9.1.7] or [18]). Note that the boundary $\partial W_{f}$ coincides with the $q_{f}$–image of the definite fold locus. As we can see easily that $\bar{f}:W_{f}\to\mathbb{R}^{2}$ is an immersion, $W_{f}$ is orientable and has nonempty boundary. Let $\alpha_{1},\alpha_{2},\ldots,\alpha_{r}$ be disjoint properly embedded arcs in $W_{f}$ such that $Y=W_{f}\setminus\left(\cup_{s=1}^{r}\mathop{\mathrm{Int}}\nolimits{N(\alpha_{s})}\right)$ is diffeomorphic to the $2$–dimensional disk (with $4r$ corner points), where $N(\alpha_{s})$ are disjoint small regular neighborhoods of $\alpha_{s}$ in $W_{f}$. We may assume that the images of the cusp points by $q_{f}$ and the crossings of $q_{f}|_{S(f)}$ lie in $\mathop{\mathrm{Int}}\nolimits{Y}$, that $q_{f}|_{S(f)}$ is transverse to each $\alpha_{s}$, and that each connected component of $q_{f}(S(f))\cap N(\alpha_{s})$ is a properly embedded arc in $N(\alpha_{s})$ intersecting with $\alpha_{s}$ exactly at one point (see the top figure of Fig. 19).

### v2

@@ 5. Maps into $\mathbb{R}^{2}$
In this section, we prove the following.
@@
Theorem 5.1 holds also for $n=2$; in fact, for each closed connected surface $M$, we can explicitly construct a generic map as described in the theorem. When $M$ is orientable, this is straightforward. For the non-orientable case, we can first construct an image simple generic map into $S^{2}$ by using [34, Theorem 1.4 (2)], for example, and then compose it with an appropriate generic map $S^{2}\to\mathbb{R}^{2}$ to get a desired generic map.
@@
There exists a generic map $f:M\to S^{2}$ which is image simple by Theorem 4.3. Let $D_{S}\cup A\cup D_{N}=S^{2}$ be a decomposition of $S^{2}$ such that $D_{S}$ and $D_{N}$ are disjoint $2$–disks, that $A$ is an annulus, that they are attached along their boundaries, and that $\mathop{\mathrm{Int}}\nolimits{D_{N}}\supset f(S(f))$. For a regular value $q\in D_{S}$, set $F=f^{-1}(q)$, which is an $(n-2)$–dimensional closed manifold. Note that $f|_{f^{-1}(D_{S}\cup A)}$ is a trivial smooth $F$–bundle over $D_{S}\cup A$ by Ehresmann’s fibration theorem [8]. In particular, we have $f^{-1}(A)\cong F\times[-1,1]\times S^{1}$.
Let $h:F\to[1,2]$ be a Morse function such that the critical values are all distinct. Define the Morse function $\varphi:F\times[-1,1]\to[1,3]$ by $\varphi(x,t)=h(x)\cos(\pi t/2)+1$, $(x,t)\in F\times[-1,1]$ (refer to [3, Fig. 40]). Then, we can construct a stable map $g:M\to\mathbb{R}^{2}$ in such a way that $g|_{f^{-1}(D_{N})}=i_{N}\circ f|_{f^{-1}(D_{N})}$, $g|_{f^{-1}(D_{S})}=i_{S}\circ f|_{f^{-1}(D_{S})}$, and $g|_{f^{-1}(A)}$ is the composition of a trivial $S^{1}$–family of Morse functions $\varphi\times\mathrm{id}_{S^{1}}:(F\times[-1,1])\times S^{1}\to[1,3]\times S^{1}$ and an appropriate embedding $[1,3]\times S^{1}\to\mathbb{R}^{2}$, where $i_{N}:D_{N}\to S^{2}$ and $i_{S}:D_{S}\to S^{2}$ are appropriate embeddings (see Fig. 13). For details, refer to [3, Proof of Theorem 7.1].
@@
Remark 5.3.
The author [22] showed that a closed orientable $3$–dimensional manifold $M$ admits an image simple generic map $f:M\to N$ without cusp points for a surface $N$ if and only if $M$ is a graph manifold. Recall that a closed orientable $3$–dimensional manifold is a graph manifold if it is the union of some $S^{1}$–bundles over compact surfaces attached along the torus boundaries. Note that the class of graph manifolds is relatively small in the set of closed orientable $3$–dimensional manifolds: for example, hyperbolic $3$–manifolds never admit such an image simple generic map.
Remark 5.4.
In this paper, basically we are studying the image of the singular point set of a generic map. For the position of the singular point set itself in the source manifold, a reader is referred to the author’s previous work [21] together with §8 of the present paper.
@@ 6. Relation to open book structures
In this section, we give an application of the above results in the odd dimensional case to open book structures.
@@
For some historical account of open book structures, the reader is referred Quinn [19, §2] and Myers [18].
@@
Let $M$ be a compact $n$–dimensional manifold with boundary. Suppose that $B$ is a proper $(n-2)$–dimensional compact submanifold of $M$ with trivial normal bundle. Here a submanifold $B$ is proper if $B\cap\partial M=\partial B$ and $B$ intersects $\partial M$ transversely. Let $\tau:N(B)\to B\times D^{2}$ be a trivialization of the normal disk bundle $N(B)$ of $B$ in $M$, where $N(B)$ is identified with a closed tubular neighborhood of $B$ in $M$ such that $N(B)\cap\partial M$ can be identified with the tubular neighborhood $N(\partial B)$ in $\partial M$. We assume that there exists a smooth $S^{1}$–bundle $\varphi:M\setminus B\to S^{1}$ such that $\varphi|_{N(B)\setminus B}$ coincides with the composition (6.1) as above and that $\varphi|_{\partial M\setminus\partial B}$ induces an open book structure with binding $\partial B$ on $\partial M$. Such a structure of $M$ is also called an open book structure; $B$ is called the binding and the closure of each fiber of $\varphi$ is called a page. Note that the restriction of the structure to $\partial M$ gives rise to an open book structure of $\partial M$ in the sense of Definition 6.1. In this case, we say that the open book structure on the boundary extends through $M$.
As an application of our results, we give a new proof to the following result originally due to Alexander [1] for $n=3$ and to Quinn [19] for $n\geq 5$ (see also [14, 18, 29, 32]).
@@
Let $bB\subset\partial M$ be the binding, $b\tau:N(bB)\to bB\times D^{2}$ the trivialization of the normal disk bundle of $bB$ in $\partial M$, and $b\varphi:\partial M\setminus bB\to S^{1}$ the fibration associated with the open book structure on $\partial M$. Let $f|_{N(bB)}:N(bB)\to\mathbb{R}^{2}$ be defined by the composition
@@
In the following, we will deform $f$ successively for our purpose and will continue to denote the resulting maps by the same symbol $f$.
We see that the number of cusps of $f$ is even, since the Euler characteristic of the pair $(M,F)$ vanishes, where $F$ is a fiber of $f$ over a point in $\partial\Delta$. Then, by the same procedure as in Levine’s cusp elimination theorem [15], we can eliminate the cusps of $f$ by pairs, since $M$ is connected. So, we may assume that $f$ has no cusps.
Now, we apply Lemma 4.8, taking into account the comments in Remark 4.9 (1) and (2), first to the definite fold image. As a result, we get a generic map $M\to\mathbb{R}^{2}$, denoted by $f$ again, such that $f(M)$ coincides with $\Delta$, the definite fold image is outward directed, and that the other fold images lie inside the interior of a $2$–disk, say $\Delta_{1}$, contained in the innermost region. Then, we apply Lemma 4.8 to $f$ restricted to $f^{-1}(\Delta_{1})$ to get a new generic map $f$ such that the fold image of absolute index $1$ is contained in $\mathop{\mathrm{Int}}\nolimits\Delta_{1}$, that it is outward directed, and that the other fold images lie inside the interior of a $2$–disk, say $\Delta_{2}$, contained in the innermost region. Repeating this procedure, we get a generic map $f:M\to\mathbb{R}^{2}$ such that $f(M)=\Delta$, that the fold images of absolute indices $i$ with $0\leq i<\lfloor(n-1)/2\rfloor$ are all directed outward, that their positions are consistent with the order of absolute indices (smaller indices in the outer position), and that the fold image of absolute index $i_{0}=\lfloor(n-1)/2\rfloor$ lies in the interior of a $2$–disk $\Delta_{i_{0}}$ contained in the innermost region. At this stage, we should note that the fold image of absolute index $i_{0}$ may not be directed outward. We may continue to assume that $f$ satisfies the conditions (1) and (2) of Definition 2.3.
@@
Now, let us first consider the case $n\geq 5$. We apply the always-realizable moves as depicted in Figs. 1417. We may assume that the disks $\Delta_{0},\Delta_{1},\ldots,\Delta_{i_{0}}$ are concentric, i.e., their centers all coincide with the origin $\mathbf{0}\in\mathbb{R}^{2}$. We denote by $\pi:\Delta\setminus\{\mathbf{0}\}\to S^{1}$ the radial projection. Let $A$ be the annulus $\Delta_{i_{0}}\setminus\mathop{\mathrm{Int}}\nolimits{\Delta_{i_{0}+1}}$, where $\Delta_{i_{0}+1}$ is a concentric disk in $\mathop{\mathrm{Int}}\nolimits{\Delta_{i_{0}}}$. We may assume that the fold image of absolute index $i_{0}$ lies in $\mathop{\mathrm{Int}}\nolimits{A}$ and that $\pi\circ f|_{S_{i_{0}}}:S_{i_{0}}\to S^{1}$ is an $S^{1}$–valued Morse function with distinct critical values, where $S_{i_{0}}$ is the set of folds of absolute index $i_{0}$ of $f$. Then, for each of the critical points, we apply the type II moves as depicted in Fig. 14 so that the images of the critical points by $f$ lie inside $\mathop{\mathrm{Int}}\nolimits{\Delta_{i_{0}+1}}$. (Before doing so, we arrange $f(S_{i_{0}})$ so that all the crossings are pushed into the regions in $A$ outside of those regions in which $f(S_{i_{0}})$ is depicted in the figure.) Note that in our case, only the top type II move (from right to left) in Fig. 5 is allowed and that the bottom type II move is not. Note also that as $n$ is odd, even if we reverse the normal orientation for a fold image of absolute index $i_{0}=(n-1)/2$, the absolute index does not change.
@@
We orient $S^{1}=\partial\Delta$ counter clockwise, for the moment, then with respect to this orientation, the numbers of local minima and maxima of the $S^{1}$–valued Morse function $\pi\circ f|_{S_{i_{0}}}$ are the same. Then, by further applying the type II moves, we may assume that the images of the critical points by $f$ lie in pairs, “face-to-face” as in the first figures of Figs. 15 and 16. Here, we can show that the fibers of $f$ over the points in $\Delta_{i_{0}+1}$ are connected by the same argument as in the proof of Lemma 4.10, since $n\geq 5$. Then, we apply always-realizable moves as depicted in Fig. 15 for each pair of the images of a local minimum and a local maximum. More precisely, we first apply type I moves to create 4 cusps. Then, we apply the cusp merge moves to the 2 pairs of cusps, which are allowed, since the fibers are connected. If we reverse the normal orientation of the newly born index–$(i_{0}+1)$ circle, then it turns into a fold image circle of absolute index $i_{0}-1$. Then, by applying the moves as depicted in Fig. 10 with $i_{0}$ replaced by $i_{0}-1$, we can turn this circle into a circle of absolute index $i_{0}$. (Note that $i_{0}-1\leq(n-3)/2$, which guarantees these moves.) Then, by applying the type II move once, we reach the final configuration in Fig. 15.
@@
Figure 15. A series of moves near a pair of critical point images
@@
Let $\Delta_{i_{0}+2}$ be a small closed $2$–disk centered at $\mathbf{0}$ disjoint from $f(S(f))$. Then, as $f$ restricted to $f^{-1}(\Delta_{i_{0}+2})$ and that to $f^{-1}(\Delta_{i_{0}+2})\cap\partial M$ are submersions, we see that $f|_{f^{-1}(\Delta_{i_{0}+2})}:f^{-1}(\Delta_{i_{0}+2})\to\Delta_{i_{0}+2}$ is a smooth fiber bundle, and it is trivial, since $\Delta_{i_{0}+2}$ is contractible. Furthermore, setting $B=f^{-1}(\mathbf{0})$, we see that $\partial B=bB$ and that the composition
@@
is a smooth fiber bundle, since $\pi\circ f$ restricted to $S(f)$ is a submersion. Hence, this gives an open book structure for $M$.
Precisely speaking, this open book structure does not induce the given open book structure on the boundary $\partial M$. However, making $N(bB)$ smaller, we see that the open book structure induced on $\partial M$ is naturally isomorphic to the given one. This means that the given open book structure on $\partial M$ extends through $M$. This completes the proof for the case $n\geq 5$.
When $n=3$, we basically follow the same procedure. However, the fibers of $f$ over the points in $\Delta_{i_{0}+1}$ may not be connected, even if the $3$–manifold $M$ is connected. Therefore, the cusp merges as in Fig. 15 may not be applicable. In this case, we can find a pair of parallel embedded curves $\gamma_{1}$ and $\gamma_{2}$ inside $M$ connecting the relevant cusp points such that their interiors avoid $S(f)$ and that their images by $f$ are immersed. Then, by imitating Alexander’s argument [1], we may further arrange $\gamma_{1}$ and $\gamma_{2}$ so that their $f$–images wind around the origin in the same direction. This is possible, since the folds have absolute indices 00 or $1$ and those of absolute index 00 are outward directed. We can, then, merge the $2$ pairs of cusps along $\gamma_{1}$ and $\gamma_{2}$, using fold-cusp crossings as well. See Fig. 18.
[…]
Figure 18. Cusp merges along curves in $M$
By choosing the parallel curves $\gamma_{1}$ and $\gamma_{2}$ appropriately, we may assume that the newly created circle of definite fold image bounds an immersed disk. We can, then, shrink the immersed disk so that we finally have an embedded one, by sliding the definite fold image circle toward the small disk embedded inside the immersed disk. Then we can safely apply the move as in Fig. 10 in order to replace the inward-directed definite fold image circle with that of absolute index $1$. Then, we can pull it toward the origin as in the last step of Fig. 15. The only difference is that now the circle may intersect with several parallel fold images of absolute index $1$, which causes no problem.
We perform these procedures for each facing pair of the images of a local maximum and a local minimum. Finally, we perform the moves as in Fig. 17. Then we get to the same conclusion as in the case of $n\geq 5$. This completes the proof. ∎
@@
(1) Note that Theorem 6.3 also covers the case with $\partial M=\emptyset$: just ignore the boundary open book structures in the proof.
(2) In the case $n=3$, we do not know if we can make the binding connected by modifying our proof.
@@ 7. Elimination of indefinite folds of certain indices
For the extended open book structure on $M$ constructed in the above proof of Theorem 6.3, the binding $B$ is connected, and the inclusion $B\to M$ induces an isomorphism $\pi_{1}(B)\cong\pi_{1}(M)$ between the fundamental groups if $n\geq 7$.
@@
Let us decompose $\Delta$, the unit disk in $\mathbb{R}^{2}$, into two half disks as $\Delta=\Delta_{+}\cup\Delta_{-}$, where $\Delta_{+}=\{(x,y)\in\Delta\,|\,y\geq 0\}$ and $\Delta_{-}=\{(x,y)\in\Delta\,|\,y\leq 0\}$. Then, we see that the inclusions $F\to f^{-1}(\Delta_{\pm})$ are homotopy equivalences. Now, the intersection $X=f^{-1}(\Delta_{+})\cap f^{-1}(\Delta_{-})$ is diffeomorphic to the union of $F$ and its copy attached along $B$ and hence we see that the inclusion $B\to X$ induces an isomorphism between the fundamental groups by the van Kampen theorem. Finally, by the van Kampen theorem applied to $M=f^{-1}(\Delta_{+})\cup_{X}f^{-1}(\Delta_{-})$, we see that the inclusion $B\to M$ induces an isomorphism between the fundamental groups. This completes the proof. ∎
@@
Let $M$ be a closed simply connected $n$–dimensional manifold with $n\geq 7$ odd. Then, there exists a stable map $f:M\to\mathbb{R}^{2}$ without cusp points such that the absolute indices $i$ of the folds satisfy either $i=0$ or $22\leq i\leq(n-1)/$ 2.
@@
As in the proof of Lemma 7.1, we consider the Morse function $g=f|_{F}:F\to[0,1]$, where $F$ is a compact $(n-1)$–dimensional manifold with boundary $\partial F=B$. By the construction of $f$, we see that the indices $i$ of the critical points satisfy $(n-1)/2\leq i\leq n-1$. On the other hand, by our assumption and Lemma 7.1, $B$ and $F$ are simply connected.
If $g$ has two or more critical points of index $n-1$, i.e. local maxima, then there exists a critical point of index $n-2$ in such a way that its dual handle connects two dual 00–handles, since $F$ is connected. Therefore, either of the two critical points of index $n-1$ and the index $n-2$ critical point make a canceling pair. (Note that by reversing the orientation of $J=[0,1]$, they make a canceling 001111 pair of critical points.) Then by [12, Part I, Chapter V, Proposition 1.1], we can cancel them leaving a pair of cusps of absolute index 00. By repeating this procedure, we may assume that $g$ has only one maximal point (refer to Fig. 19).
@@
Let $U$ be a small open regular neighborhood of $J$ in $\mathbb{R}^{2}$. Note that $\Delta\setminus U$ is naturally identified with $[0,1]\times[-1,1]$, where the first factor corresponds to a radial outward arc and the second factor corresponds to the direction of $\partial\Delta=S^{1}$. Then, $f$ restricted to $f^{-1}(\Delta\setminus U)$ is considered to be a smooth $1$–parameter family of functions from $F$ to $[0,1]$ parametrized by $[-1,1]$. Now, by applying an argument similar to that in the proof of [12, Part I, Chapter V, Proposition 3.5], we can homotopically modify $f$ so that it has only one fold circle of absolute index 00 and that it is embedded into $\mathbb{R}^{2}$ in the outermost place with the expense of creating some fold circles of absolute index $2$.
Since $F$ is simply connected of dimension greater than or equal to $6$, by [17, Theorem 8.1], we can cancel each critical point of index $1$ with the expense of creating critical points of index $3$. Then, again by the same argument as in the proof of [12, Part I, Chapter V, Proposition 3.5], we can homotopically modify $f$ so that it has no fold points of absolute index $1$ and that it has only one fold circle of absolute index 00 embedded into $\mathbb{R}^{2}$ in the outermost place. This completes the proof. ∎
Now, as a small digression, let us consider the topology of manifolds admitting a generic map into $\mathbb{R}^{2}$ without fold of absolute index $1$.
@@
We may assume that $f$ is a stable map. Then, as $f$ does not have folds of absolute index $1$, we can show that $W_{f}$ has the structure of a compact connected smooth $2$–dimensional manifold (for example, see [33, Theorem 9.1.7] or [25]). Note that $q_{f}$ restricted to the definite fold locus is a diffeomorphism onto the boundary $\partial W_{f}$. As we can see easily that $\bar{f}:W_{f}\to\mathbb{R}^{2}$ is an immersion, $W_{f}$ is orientable and has nonempty boundary. Let $\alpha_{1},\alpha_{2},\ldots,\alpha_{r}$ be disjoint properly embedded arcs in $W_{f}$ such that $Y=W_{f}\setminus\left(\cup_{s=1}^{r}\mathop{\mathrm{Int}}\nolimits{N(\alpha_{s})}\right)$ is diffeomorphic to the $2$–dimensional disk (with $4r$ corner points), where $N(\alpha_{s})$ are disjoint small regular neighborhoods of $\alpha_{s}$ in $W_{f}$. We may assume that the images of the cusp points by $q_{f}$ and the crossings of $q_{f}|_{S(f)}$ lie in $\mathop{\mathrm{Int}}\nolimits{Y}$, that $q_{f}|_{S(f)}$ is transverse to each $\alpha_{s}$, and that each connected component of $q_{f}(S(f))\cap N(\alpha_{s})$ is a properly embedded arc in $N(\alpha_{s})$ intersecting with $\alpha_{s}$ at exactly one point (see the top figure of Fig. 20).